The proublem is:

The school production of our school was a big success. For opening night, 410 tickets were sold. Students paid $3 each, while non-students paid $7 each. If a total of $1650 was collected, how many students and how many non-students attended?

I tried using an algebra equation to answer the question but I couldn't figure it out. The answer I got was 123= students and 287= non-students. Although I did try guess and check and I came up with the correct answer 105= students and 305= non-students. Can you please help me come up with an equation?

Answer by Paul(988)

students = x
non-students=y
x+y=410
y=410-x
3x+7y=1650
3x+7(410-x)=1650
3x+2870-7x=1650
-4x=-1220 (divide both sides by -4)
x=305 y=410-305=105
Hence, 305 students attended and 105 non-students attended